OpenStudy (anonymous):

Check my answer? Simplify: 4i^4 -3i^3 -sqrt-3 sqrt3 A. 3-3i^3 +4i^4 B. 7+3i C. 4-6i D. 4i E. none of these I'm getting D. 4i ?

OpenStudy (anonymous):

@Nnesha

OpenStudy (danjs):

This is the property \[i^2 = -1~~~or~~~i=\sqrt{-1}\]

OpenStudy (danjs):

i = i i^2 = -1 i^3 = i^2 *i = -1*i i^4 = i^2 *i^2 = -1 * -1 = 1 The cycle repeats i -1 -i +1

OpenStudy (danjs):

\[4i^4-3i^3-\sqrt{-3}\sqrt{3}\] Did i get that right?

OpenStudy (danjs):

\[4 * i^2*i^2 - 3*i^2*i - \sqrt{-1}\sqrt{3}\sqrt{3}\] replace all i^2 by -1 , all root -1 by i

OpenStudy (danjs):

rot -1 is i^2

OpenStudy (anonymous):

Okay so far it looks like how i did it.

OpenStudy (danjs):

4*-1*-1 - 3*-1 *i - i^2 *3 4 + 3i -3i^2 4 + 3i - 3*-1 4 + 3i + 3 7 + 3i

OpenStudy (danjs):

you see each part?

OpenStudy (anonymous):

yess, i see each part.

OpenStudy (anonymous):

i think i went wrong on the 3rd step somehow

OpenStudy (danjs):

just treat these like normal probs with the 'i' being a normal variable, then just use the property that i^2 = -1 or square root of -1 = i

OpenStudy (anonymous):

okay ill write that tip down. thank you :)

OpenStudy (danjs):

welcome